# Toward Intelligent and Secure Cloud: Large Language Model Empowered Proactive Defense

### v1

@@ Abstract
The rapid evolution of cloud computing technologies and the increasing number of cloud applications have provided a large number of benefits in daily lives. However, the diversity and complexity of different components pose a significant challenge to cloud security, especially when dealing with sophisticated and advanced cyberattacks. Recent advancements in generative foundation models (GFMs), particularly in the large language models (LLMs), offer promising solutions for security intelligence. By exploiting the powerful abilities in language understanding, data analysis, task inference, action planning, and code generation, we present LLM-PD, a novel proactive defense architecture that defeats various threats in a proactive manner. LLM-PD can efficiently make a decision through comprehensive data analysis and sequential reasoning, as well as dynamically creating and deploying actionable defense mechanisms on the target cloud. Furthermore, it can flexibly self-evolve based on experience learned from previous interactions and adapt to new attack scenarios without additional training. The experimental results demonstrate its remarkable ability in terms of defense effectiveness and efficiency, particularly highlighting an outstanding success rate when compared with other existing methods.
@@ I Introduction
The rapid technological advancement of cloud computing has notably accelerated recently, establishing itself as a significant paradigm in service-oriented computing. In the last decade, cloud computing has received huge attention due to its notable advantages such as scalability, flexibility, and increased cost-efficiency. For example, cloud computing has entangled with the daily lives of businesses and individuals, offering a diverse range of services (e.g., storage, networking, and platform provisions) seamlessly accessible from any geographical location connected to the Internet.
Despite the obvious benefits of cloud computing, the diverse and complex nature of its various components, i.e., network, architecture, APIs, and hardware, has led to heightened security apprehensions. Due to the utilization of standard Internet protocols and virtualization methods, it is susceptible to potential security breaches. These vulnerabilities may stem from conventional sources such as IP spoofing, Address Resolution Protocol, and Distributed Denial of Service (DDoS) attacks, among others. Additionally, emerging threats such as zero-day attacks, characterized by their unknown nature, pose a considerable challenge to traditional solutions, which may not be sufficiently effective in mitigating such attacks.
@@
Although these solutions overcome, to some extent, the shortcomings of traditional solutions, they require modification of the mitigation mechanisms which may fail across diverse environments. Moreover, the decision-making of defense deployment predominantly relies on heuristic, machine learning, and deep learning algorithms. Nevertheless, given the increasing complexity of cloud-based applications and the wide array of attack vectors, it is hard for any specific strategy to fully adapt to the different and time-varying scenarios in the cloud. For this reason, there is an urgent requirement for an intelligent and adaptable guidance to facilitate proactive defense within the cloud environment.
Fortunately, as a typical example of Generative Foundation Models (GFMs) that can be adapted for specific purposes, Large Language Models (LLMs) have made a profound impact on research and practical applications [4]. Specially, LLMs excel in addressing general tasks through in-context learning, generating multimodal data via prompt engineering, breaking down complex problems into simpler components through strategic reasoning, simulating specified scenarios through role-playing, and making decisions based on prior knowledge and situational understanding [5]. Building upon these advantages, recent research has leveraged LLMs in the cybersecurity community, empowering security professionals to explore various attack vectors (e.g., vulnerability detection) and develop autonomous agents (e.g., code fixing). Therefore, the extensive knowledge encoded in LLMs has sparked our interest in explore their potential for comprehensive protection in the complex and ever-changing cloud security scenarios.
@@
  • We design a novel robust and efficient defense architecture called LLM-PD, with essential components of collection, assessment, decision, deployment, and feedback, for improved intelligent and secure cloud. To the best of our knowledge, this is the first solution that integrates the ability of LLMs to protect cloud networks proactively from advanced evolving attacks and cyber-threats.
  • We conduct a detailed case study utilizing GPT-4o mini, a pre-trained API-based LLM from OpenAI, to demonstrate how our architecture can be practically implemented and how it automatically adapts to different types of DoS attacks over the cloud environment. The source code is publicly available at https://github.com/SEU-ProactiveSecurity-Group/LLM-PD.
  • We evaluate the performance of LLM-PD and perform comparative analyses with other existing approaches. Experimental results highlight the advantages of this architecture in terms of surviving rate, time efficiency, and success rate across various attack scenarios. Noteworthy future research directions are also discussed.
@@ II-A Proactive Defense in Cloud Networks
Proactive defense empowers cybersecurity personnel with anticipatory control over system defenses, establishing an advantage and enhancing protection to a more sophisticated level. Different from the reactive approach, it involves real-time monitoring of the system, enabling the rapid detection of changes in network traffic, analysis of program behavior, and the prevention of suspicious activities to ensure security.
There have been some efforts that attempted to enhance security in a proactive manner. For example, Zhou et. al [6] combine MTD technologies and Deep Reinforcement Learning (DRL) in a seamless way to proactively defend against Low-rate DDoS attacks. This approach introduces a Markov decision process (MDP) to formalize attack scenarios and adopts a Deep Q-network (DQN) algorithm to mitigate the impact of attacks in a resource-efficient manner.
Ma et. al [7] leverage the mutation-enabled defense mechanism in the cloud-native applications for mitigating Man-in-The-Middle (MITM) attacks. Driven by a maximizing coloring-based mutation grouping algorithm, it improves the robustness against attacks while guaranteeing the continuity of network connection with an acceptable performance loss.
Wu et. al [8] propose a security defense framework based on Intrinsic Cloud Security (ICS). It fuses the MTD concept and mimic defense theory within the Network Functions Virtualization (NFV) based cloud and subsequently implements the integration to achieve proactive protection against multiple threats, including cache side-channel attacks, code reuse attacks, and latent persistent threats.
However, existing work exhibits certain shortcomings, notably a lack of intelligent and adaptive strategic guidance, which is crucial for preventing advanced and sophisticated intruders. Although some methods incorporate intelligent algorithms, they often require extensive training over thousands of episodes. Moreover, the policies learned in one specific configuration typically do not generalize well and necessitate retraining when applied to even slightly different scenarios.
@@ II-B LLM for Cybersecurity Enhancement
Traditional approaches, such as signature-based detection and rule-based systems, often struggle to keep pace with the ever-changing threat landscape in the dynamic cyberspace. However, advancements in LLMs have marked in a significant shift in cybersecurity paradigms, offering more adaptive and intelligent technologies [9]. Specifically, functionalities of LLMs have contributed to promising enhancements in defense applications, including threat detection, vulnerability analysis, automated defense mechanisms, and others.
@@
Furthermore, Yin et. al [11] propose a pipeline for quantitatively evaluating LLMs’ performance in various software vulnerability-related tasks using publicly available datasets. Their evaluation results indicate that while LLMs perform well in certain aspects, there is still a need for improvement in understanding the subtle differences in code vulnerabilities and in their ability to describe these vulnerabilities, in order to fully realize their potential.
@@
One of the main issues with these works is that the comprehensive aspect of the mitigation problem, which is not limited to the implementation of a single security feature, is often missing. Defending against complex threat vectors entails lifecycle protection over a relatively long time period and necessitates effective responses to attacker behaviors.
@@ III LLM Empowered Proactive Defense Architecture for Cloud Security
To achieve a self-autonomous, AI-driven, and secure-by-design defense implementation, we present an architecture, namely LLM-PD, that deeply integrates LLMs with existing security features. This integration aims to deliver intelligent, efficient, scalable, and proactive security enhancement across cloud networks. Fig. 1 illustrates the architecture of LLM-PD, which consists of five essential components as below.
@@ III-A Data Collection and Reconstruction
In cloud networks, the extensive deployment of network applications results in a substantial volume of data available for security purposes. This data includes network traffic, performance metrics, logs, events, and service configurations, etc. Essentially, each type of data is considered as independent and is formatted differently. For example, Wireshark is a tool that used to capture network traffic, typically in .pcap format, whereas alarm events are generated in .txt format based on firewall rules. Therefore, the ability to proficiently use multiple security tools and analyze massive heterogeneous data poses a significant challenge, even for cybersecurity professionals.
Powered by the extraordinary prowess of LLMs, this component works as a collector, which is generally fed with a prompt of natural language descriptions that cover the range of targets, tools, time, and reconstruction format. In detail, the collector is first designed to call multiple security tools to gather data from the target cloud network within given time period. Specially, when new collection tools are introduced in the future, this collector can adapt by modifying its prompts easily. Next, to extract key concepts from multiple forms of data, it requires converting the data into text format and obtaining a brief report on the cloud network. Nevertheless, given the large number of data sources, redundant and irrelevant information is inevitable. To this end, the collector aggregates data by evaluating the relationships between them and eliminating duplicate contents. Finally, this module reconstructs the refined results into a standardized representation (e.g., JSON file) and transmits it to the next stage for further processing.
@@ III-B Status and Risk Assessment
Upon receiving the formatted information from the collector, we necessitate another module prompted to serve as an analyzer to perform a thorough assessment of the data. To effectively ascertain the system’s status and identify potential risks, this module encompasses two primary functions.
@@ III-B2 Risk Evaluation
The risk evaluation function primarily focuses on the security indicators of cloud networks. This function utilizes information collected from various monitoring tools, in conjunction with historical normal data, to identify potential threats and anomalies within the system. Concurrently, the risk level is quantified (e.g., on a scale of 0-10) based on the scope, impact, and duration of the threat, thereby prioritizing following defense tasks. Additionally, in cases of cloud network failures caused by severe attacks or malicious disruptions (such as service interruptions), this feature can also leverage existing information for fault localization, effectively identifying attack targets and generating defense goals.
Finally, the analyzer integrates the results of both functions and output them to the subsequent module. It then awaits the arrival of new data to initiate the next round of analysis.
@@ III-C1 Task Decomposition
In a high-level threat scenario involving multiple defense objectives, it is significant to clearly plan and generate detailed tasks for better strategic reasoning. To this end, we have designed a task decomposition function that breaks down complex tasks into multiple independent subtasks. We require the LLM to delineate the implementation constraints of each subtask, assign priorities based on the previously mentioned risk levels, and analyze the dependencies among tasks for sequential arrangement.
@@ III-C2 Inference and Decision-Making
In this process, tasks are executed sequentially according to their assigned order. For each subtask, this function systematically deduces the necessary step-by-step process required for decision-making, with the logic outlined as follows. First, details of the task are extracted from each subtask. Then, based on the current task requirements and constraints, human-like sequential reasoning for defensive actions over future attack-defense interactions is conducted. Additionally, this function is designed not only to enhance security but also to incorporate preferences, such as focusing on reducing defense overhead or ensuring user experience. These preferences can be preset by security personnel and dynamically adjusted during implementation. Finally, by balancing defense effectiveness and preferences, it imports solvers to optimize the policy among the available strategies.
@@ III-D Defense Deployment and Execution
When a particular mechanism does not exist in the defense library, the deployer harnesses the capabilities of LLMs to invoke programming languages, thereby generating useful scripts based on carefully crafted prompts to construct the necessary defense mechanism. Once these scripts are generated, they undergo a compilation process to ensure their operational integrity and compatibility with the cloud environment. Upon successful compilation, the scripts are deployed and executed within the cloud infrastructure, effectively implementing the desired defense mechanism according to the generated policy. Simultaneously, these newly generated scripts are archived in the defense library, enriching the repository with new capabilities and ensuring that similar requirements in the future can be addressed more efficiently.
This module not only enhances the adaptability and resilience of the defense system but also ensures that the deployment is both efficient and effective. Through this meticulous approach empowered by LLMs, the deployer allows itself to evolve dynamically in response to providing comprehensive protection against emerging threats.
@@ III-E Effectiveness Analysis and Feedback
Specifically, the evaluation of defense effectiveness is not limited to the analysis of a single factor. In this framework, we incorporate additional parameters beyond security, such as recovery time, resource consumption, financial cost, and Quality of Service (QoS), to achieve a more comprehensive and rational assessment of effectiveness. Among these parameters, security remains the core element, signifying the capability to effectively mitigate cloud threats. Recovery time, on the other hand, is a critical measure of efficiency, reflecting the system’s ability to swiftly restore cloud services to their normal operational state following an incident. Resource expenditure and financial cost provide insights into the expenses incurred by the defense actions. Meanwhile, QoS is assessed from the user perspective, highlighting the potential impact of defense actions on application performance and user experience.
Subsequently, we introduced a validation function to rigorously assess the effectiveness of the defense measures. This function serves a dual purpose: initially, it evaluates the efficacy of the strategies generated during the defense process, ensuring that they meet the desired objectives and effectively mitigate identified threats. Secondly, it verifies the correct implementation of the defense actions, confirming that they are executed as intended and align with the strategic goals. Additionally, this function annotates the current round’s strategies and actions based on the validation results and transmits these annotations to the memory module for storage. The memory module plays a pivotal role in the self-improvement and evolution of the defense architecture by feeding both positive and negative memories back to other LLM agents. This iterative feedback loop enables a dynamic learning process and allows the system to adapt and refine its strategies over time, enhancing its resilience and effectiveness in the face of evolving security challenges. Importantly, this entire process occurs without the need for human intervention or expert knowledge, highlighting the autonomous and robust nature of the proposed protection.

### v2

@@ Abstract
The rapid evolution of cloud computing technologies and the increasing number of cloud applications have provided numerous benefits in our daily lives. However, the diversity and complexity of different components pose a significant challenge to cloud security, especially when dealing with sophisticated and advanced cyberattacks. Recent advancements in generative foundation models (GFMs), particularly in the large language models (LLMs), offer promising solutions for security intelligence. By exploiting the powerful capabilities in language understanding, data analysis, task inference, action planning, and code generation, we present LLM-PD, a novel defense architecture that proactively defeats various threats. LLM-PD can efficiently make decisions through comprehensive data analysis and sequential reasoning, as well as dynamically create and deploy actionable defense mechanisms on the target cloud. Furthermore, it can flexibly self-evolve based on experience learned from previous interactions and adapt to new attack scenarios without additional training. The experimental results demonstrate its remarkable ability in terms of defense effectiveness and efficiency, particularly highlighting an outstanding success rate when compared with other existing methods.
@@ I Introduction
Cloud computing has accelerated rapidly in recent years, establishing itself as a significant paradigm in service-oriented computing. In the last decade, cloud computing has received significant attention due to its notable advantages such as scalability, flexibility, and increased cost-efficiency. For example, cloud computing has been integrated into the daily lives of businesses and individuals, offering a diverse range of services (e.g., storage, networking, and platform provisions) seamlessly accessible from any geographical location connected to the Internet.
Despite the obvious benefits of cloud computing, the diverse and complex nature of its various components, i.e., network, architecture, APIs, and hardware, has led to heightened security apprehensions. Due to the utilization of standard Internet protocols and virtualization methods, it is susceptible to potential security breaches. These vulnerabilities may stem from conventional sources such as IP spoofing, Address Resolution Protocol spoofing, and Distributed Denial of Service (DDoS) attacks, among others. Additionally, emerging threats such as zero-day attacks, characterized by their unknown nature, pose a considerable challenge to traditional solutions, which may not be sufficiently effective in mitigating such attacks.
@@
Although these solutions overcome, to some extent, the shortcomings of traditional solutions, they require modifications to mitigation mechanisms that may not be effective across diverse environments. Moreover, the decision-making of defense deployment predominantly relies on heuristic, machine learning, and deep learning algorithms. Nevertheless, given the increasing complexity of cloud-based applications and the wide array of attack vectors, it is hard for any specific strategy to fully adapt to the different and time-varying scenarios in the cloud. For this reason, there is an urgent requirement for an intelligent and adaptable approach to facilitate proactive defense within the cloud environment.
Fortunately, as a typical example of Generative Foundation Models (GFMs) that can be adapted for specific purposes, Large Language Models (LLMs) have made a profound impact on research and practical applications [4]. Specifically, LLMs excel in addressing general tasks through in-context learning, generating multimodal data via prompt engineering, breaking down complex problems into simpler components through strategic reasoning, simulating specified scenarios through role-playing, and making decisions based on prior knowledge and situational understanding [5]. Building upon these advantages, recent research has leveraged LLMs in the cybersecurity community, empowering security professionals to explore various attack vectors (e.g., vulnerability detection) and develop autonomous agents (e.g., code fixing). Therefore, the extensive knowledge encoded in LLMs has sparked our interest in exploring their potential for comprehensive protection in the complex and ever-changing cloud security scenarios.
@@
  • We design a novel robust and efficient cloud security architecture called LLM-PD. To the best of our knowledge, this is the first comprehensive end-to-end LLM-driven proactive defense architecture where specialized agents handle different aspects of the defense lifecycle while maintaining contextual awareness.
  • We achieve cross-attack adaptability without extensive training, overcoming limitations of traditional ML methods that require retraining for different scenarios. Moreover, we develop a self-evolving memory feedback mechanism that enables continuous learning and improvement of defense capabilities through experience.
  • We conduct a case study with GPT-4o mini to demonstrate practical implementation of our architecture and its adaptation to various DoS attacks. Comparative experiments reveal significant advantages in surviving rate, time efficiency, and success rate over existing approaches. The source code is available at https://github.com/SEU-ProactiveSecurity-Group/LLM-PD.
@@ II-A Proactive Defense in Cloud Networks
Proactive defense empowers cybersecurity personnel with anticipatory control over system defenses, establishing a tactical advantage and elevating protection to a more sophisticated level. Unlike the reactive approach, it involves real-time monitoring of the system, enabling the rapid detection of changes in network traffic, analysis of program behavior, and the prevention of suspicious activities to ensure security.
There have been some efforts that attempted to enhance security in a proactive manner. For example, Zhou et al. [6] seamlessly combine MTD technologies and Deep Reinforcement Learning (DRL) to proactively defend against Low-rate DDoS attacks. This approach introduces a Markov decision process (MDP) to formalize attack scenarios and adopts a Deep Q-network (DQN) algorithm to mitigate the impact of attacks in a resource-efficient manner.
Ma et al. [7] leverage the mutation-enabled defense mechanism in the cloud-native applications for mitigating Man-in-The-Middle (MITM) attacks. Driven by a maximizing coloring-based mutation grouping algorithm, it improves the robustness against attacks while guaranteeing the continuity of network connection with an acceptable performance loss.
Wu et al. [8] propose a security defense framework based on Intrinsic Cloud Security (ICS). It fuses the MTD concept and mimic defense theory within the Network Functions Virtualization (NFV) based cloud and subsequently implements the integration to achieve proactive protection against multiple threats, including cache side-channel attacks, code reuse attacks, and latent persistent threats.
However, existing work lacks intelligent decision-making and typically relies on predefined rules or limited policy sets that cannot adapt to complex and evolving threats. Although some methods incorporate intelligent algorithms, they require specialized training over thousands of episodes for each attack type and do not generalize well to even slightly different scenarios without extensive retraining. In contrast, LLM-PD offers cross-scenario defense capabilities without requiring retraining for different attack scenarios, and can dynamically generate and deploy entirely new defense measures.
@@ II-B LLM for Cybersecurity Enhancement
Traditional approaches, such as signature-based detection and rule-based systems, often struggle to keep pace with the ever-changing threat landscape in the dynamic cyberspace. However, advancements in LLMs have marked a significant shift in cybersecurity paradigms, offering more adaptive and intelligent technologies [9]. Specifically, the functionalities of LLMs have contributed to promising enhancements in defense applications, including threat detection, vulnerability analysis, automated defense mechanisms, and others.
@@
Furthermore, Yin et al. [11] propose a pipeline for quantitatively evaluating LLMs’ performance in various software vulnerability-related tasks using publicly available datasets. Their evaluation results indicate that while LLMs perform well in certain aspects, there is still a need for improvement in understanding the subtle differences in code vulnerabilities and in their ability to describe these vulnerabilities, in order to fully realize their potential.
@@
Existing LLM applications in security primarily focus on isolated security functions and cannot ensure lifecycle protection over a relatively long time period. On the contrary, we integrate complete defense lifecycle into a cohesive end-to-end intelligent system and realize continuous self-optimization through memory-based iterative learning.
@@ III LLM Empowered Proactive Defense Architecture for Cloud Security
To achieve an autonomous, AI-driven, and secure-by-design defense implementation, we present an architecture, named LLM-PD, that deeply integrates LLMs with existing security features, as illustrated in Fig. 1. LLM-PD goes beyond simple prompt engineering by introducing a sophisticated, security-specific multi-agent system designed for cloud environments. Each component incorporates a specific LLM-driven agent, all of which together build up a comprehensive proactive defense pipeline.
@@ III-A Data Collection and Reconstruction
In cloud networks, the extensive deployment of network applications results in a substantial volume of data available for security purposes. This data includes network traffic, performance metrics, logs, events, and service configurations, etc. Essentially, each type of data is considered independent and is formatted differently. For example, Wireshark is a tool that used to capture network traffic, typically in .pcap format, whereas alarm events are generated in .txt format based on firewall rules. Therefore, the ability to proficiently use multiple security tools and analyze massive heterogeneous data poses a significant challenge, even for cybersecurity professionals.
Powered by the extraordinary prowess of LLMs, this component works as a collector, which is generally fed with a prompt of natural language descriptions that cover the range of targets, tools, time, and reconstruction format. In detail, the collector is first designed to call multiple security tools to gather data from the target cloud network within given time period. Specifically, this collector can adapt to new collection tools through natural language interactions. Different from typical approaches, our collector implements semantically-driven heterogeneous data integration that can process unstructured security data from multiple sources. Nevertheless, given the large number of data sources, redundant and irrelevant information is inevitable. To this end, the collector aggregates data by evaluating the relationships between them and eliminating duplicate contents. Finally, this module reconstructs the refined results into a standardized representation (e.g., JSON file) and transmits it to the next stage for further processing.
@@ III-B Status and Risk Assessment
Upon receiving the formatted information from the collector, we necessitate another module to serve as an analyzer to perform a thorough assessment of system status and identify potential security risks.
@@ III-B2 Risk Evaluation
The risk evaluation function primarily focuses on the security indicators of cloud networks. This function utilizes information collected from various monitoring tools, in conjunction with historical normal data, to identify potential threats and anomalies within the system. Concurrently, the risk level is quantified (e.g., on a scale of 0-10) based on the scope, impact, and duration of the threat, thereby prioritizing subsequent defense tasks. Additionally, in cases of cloud network failures caused by severe attacks or malicious disruptions (such as service interruptions), this feature can also leverage existing information for fault localization, effectively identifying attack targets and generating defense goals.
Unlike rule-based analyzers, we implement multi-level threat modeling and quantitative risk assessment combined with causal reasoning to understand the relationships between different attack indicators and system anomalies.
@@ III-C1 Task Decomposition
In a high-level threat scenario involving multiple defense objectives, it is crucial to clearly plan and generate detailed tasks for better strategic reasoning. To this end, we have designed a task decomposition function that breaks down complex tasks into multiple independent subtasks. We require the LLM to delineate the implementation constraints of each subtask, assign priorities based on the previously mentioned risk levels, and analyze the dependencies among tasks for sequential arrangement.
@@ III-C2 Inference and Decision-Making
In this process, tasks are executed sequentially according to their assigned order. For each subtask, this function systematically deduces the necessary step-by-step process required for decision-making, with the logic outlined as follows. First, details of the task are extracted from each subtask. Then, based on the current task requirements and constraints, human-like sequential reasoning for defensive actions over future attack-defense interactions is conducted. Additionally, this function is designed not only to enhance security but also to incorporate preferences, such as focusing on reducing defense overhead or ensuring user experience. These preferences can be preset by security personnel and dynamically adjusted during implementation. Finally, by balancing multiple competing objectives, it imports solvers to optimize the policy among the available strategies.
@@ III-D Defense Deployment and Execution
When a particular mechanism does not exist in the defense library, the deployer harnesses the capabilities of LLMs to invoke programming languages, thereby generating useful scripts based on carefully crafted prompts to construct the necessary defense mechanism. Once these scripts are generated, they undergo a compilation and verification process to ensure their reliability and operational compatibility, addressing a critical limitation in automated defense generation that distinguishes our approach from existing work. Upon successful verification, the scripts are deployed and executed within the cloud infrastructure, effectively implementing the desired defense mechanism according to the generated policy. Simultaneously, these newly generated scripts are archived in the defense library, enriching the repository with new capabilities and ensuring that similar requirements in the future can be addressed more efficiently.
@@ III-E Effectiveness Analysis and Feedback
III-E1 Multi-dimensional Evaluation
The evaluation of defense effectiveness in LLM-PD employs a multi-dimensional assessment approach that goes beyond analyzing a single security metric. Our framework integrates security effectiveness, recovery time, resource consumption, financial cost, and Quality of Service (QoS) to achieve a more comprehensive and rational assessment of effectiveness. Each dimension is assigned an importance weight based on the specific defense context and organizational priorities. This quantitative effectiveness score enables the LLM agent to make nuanced decisions that balance multiple competing objectives.
III-E2 Endpoint Validation
The primary purpose of the endpoint validation function is to determine whether the current defense process has reached its endpoint. After each defensive action is deployed, this function evaluates the current system state against success mitigation and failure conditions. When the validation function determines that the defense round has concluded, it tags the complete sequence of defensive actions implemented during the current episode with outcome flags (success or failure) and relevant contextual metadata. The annotated defense sequence is then transmitted to the memory module for storage and future reference.
III-E3 Memory-Based Optimization
The memory module implements a sophisticated dual-layer memory architecture that consists of intra-episode pool and inter-episode pool. Within each defense episode, the intra-episode memory pool maintains action-effectiveness sequences that associate each defensive action with quantitative effectiveness metrics. These action-effectiveness pairs guide the LLM agent toward selecting progressively higher-effectiveness defensive actions within the ongoing episode. For the inter-episode memory pool, it implements a sliding window mechanism that stores the most recent batches of validation results, that is, complete defense action sequences along with their overall outcomes. This module addresses token limit constraints while preserving the most relevant historical knowledge, prioritizing successful action sequences that required the fewest steps and avoiding previously failed defense actions. This iterative feedback loop enables a dynamic learning process and allows the system to adapt and refine its strategies over time, enhancing its resilience and effectiveness in the face of evolving security challenges. Importantly, this entire process occurs without the need for human intervention or expert knowledge, facilitating continuous improvement across multiple defense episodes.
III-F Practical Applications and Deployment Scenarios
The LLM-PD architecture can be effectively deployed across various cloud security scenarios. We highlight three critical application areas as follows.
III-F1 Mitigating DoS/DDoS Attacks in Cloud Services
LLM-PD protects against various DoS attacks, such as SYN flooding, SlowHTTP, and Memory DoS. The architecture dynamically adjusts security parameters, implements rate limiting, and executes appropriate mitigation strategies based on attack characteristics, ensuring continuous service availability.
III-F2 Enterprise Cloud Security Operations
In enterprise networks, LLM-PD enhances security operations by monitoring cloud infrastructure, automating alert triage, and implementing initial containment actions before human involvement. This significantly reduces detection and response times.
III-F3 Multi-Cloud Defense Orchestration
For organizations employing multi-cloud strategies, LLM-PD serves as a unified defense orchestrator that normalizes security telemetry across diverse cloud platforms. This enables coordinated defense actions across provider boundaries, preventing attackers from exploiting differences between cloud environments.